Solve the following Non-Homogeneous systems by Gauss elimination and Gauss




Jordan Method.




a) 2x + y + 4z = 12




8x - 3y + 2z = 20 4x + 11y - z = 33

1
Expert's answer
2021-12-28T11:18:02-0500

Solution:

By Gauss elimination (or Gauss jordan):

"\left\{\begin{array}{l}

2 x+y+4 z=12 \\

8 x-3 y+2 z=20 \\

4 x+11 y-z=33

\end{array}\right."

Rewrite the system in matrix form and solve it by Gaussian Elimination (Gauss-Jordan elimination)

"\left(\begin{array}{ccc|c}

2 & 1 & 4 & 12 \\

8 & -3 & 2 & 20 \\

4 & 11 & -1 & 33

\end{array}\right)"

"\mathrm{R}_{1} / 2 \rightarrow \mathrm{R}_{1}" (divide the 1 row by 2 )

"\left(\begin{array}{ccc|c}

1 & 0.5 & 2 & 6 \\

8 & -3 & 2 & 20 \\

4 & 11 & -1 & 33

\end{array}\right)"

"\mathrm{R}_{2}-8 \mathrm{R}_{1} \rightarrow \mathrm{R}_{2}" (multiply 1 row by 8 and subtract it from 2 row);

"\mathrm{R}_{3}-4 \mathrm{R}_{1} \rightarrow \mathrm{R}_{3}" (multiply 1 row by 4 and subtract it from 3 row)

"\left(\begin{array}{ccc|c}

1 & 0.5 & 2 & 6 \\

0 & -7 & -14 & -28 \\

0 & 9 & -9 & 9

\end{array}\right)"

"\mathrm{R}_{2} /-7 \rightarrow \mathrm{R}_{2}" (divide the 2 row by -7 )

"\left(\begin{array}{ccc|c}

1 & 0.5 & 2 & 6 \\

0 & 1 & 2 & 4 \\

0 & 9 & -9 & 9

\end{array}\right)"

"\mathrm{R}_{1}-0.5 \mathrm{R}_{2} \rightarrow \mathrm{R}_{1}" (multiply 2 row by 0.5 and subtract it from 1 row);

"\mathrm{R}_{3}-9 \mathrm{R}_{2} \rightarrow \mathrm{R}_{3}" (multiply 2 row by 9 and subtract it from 3 row)

"\left(\begin{array}{ccc|c}

1 & 0 & 1 & 4 \\

0 & 1 & 2 & 4 \\

0 & 0 & -27 & -27

\end{array}\right)"

"\mathrm{R}_{3} /-27 \rightarrow \mathrm{R}_{3}" (divide the 3 row by -27 )

"\left(\begin{array}{lll|l}

1 & 0 & 1 & 4 \\

0 & 1 & 2 & 4 \\

0 & 0 & 1 & 1

\end{array}\right)"

"\mathrm{R}_{1}-1 \mathrm{R}_{3} \rightarrow \mathrm{R}_{1}" (multiply 3 row by 1 and subtract it from 1 row);

"\mathrm{R}_{2}-2 \mathrm{R}_{3} \rightarrow \mathrm{R}_{2}" (multiply 3 row by 2 and subtract it from 2 row)

"\left(\begin{array}{lll|l}

1 & 0 & 0 & 3 \\

0 & 1 & 0 & 2 \\

0 & 0 & 1 & 1

\end{array}\right)"

Thus, "x=3,y=2,z=1"